OpenStudy (anonymous):

A particle is moving with velocity v(t) = t^2 – 9t + 18 with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero, that is, s(0) = 1. 1:The average velocity over the interval 0 to 8 seconds 2:The time interval(s) when the particle is moving right 3:The time interval(s) when the particle is a: going faster b: slowing down

OpenStudy (anonymous):

The mean value theorem for integrals will help with part (1). \[\text{average velocity}=\frac{1}{8-0}\int_0^8v(t)\,dt\] Identically, you can solve for the position function \(s(t)\), given by \(s(t)=\int v(t)\,dt\) and use the initial condition. Then the mean value theorem says that the average rate of change of the position (i.e. average velocity) over the time interval would be \[\text{average velocity}=\frac{s(8)-s(0)}{8-0}\] You should notice that the numerators for both expressions are the same as a consequence of the fundamental theorem of calculus.

OpenStudy (anonymous):

@ can you show me the steps of how it done? i really don't know what is going on.

OpenStudy (anonymous):

What particularly are you having trouble with? Integrating? Recall the power rule: \[\int x^n\,dx=\dfrac{1}{n+1}x^{n+1}+C\]for \(n\neq-1\). Here you have \[\frac{1}{8}\int_0^8(t^2-9t+18)\,dt=\frac{1}{8}\left[\frac{1}{3}t^3-\frac{9}{2}t^2+18t\right]_0^8=\cdots\] ================================================== Or, using the second approach, you solve for \(s(t)\): \[\begin{align*}s(t)&=\int v(t)\,dt\\\\ &=\frac{1}{3}t^3-\frac{9}{2}t^2+18t+C \end{align*}\]When \(t=0\), you have \(s(0)=1\), so \[\begin{align*}s(0)&=\frac{1}{3}0^3-\frac{9}{2}0^2+18(0)+C\\\\ 1&=0-0+0+C\end{align*}\]So the position function is given by \[s(t)=\frac{1}{3}t^3-\frac{9}{2}t^2+18t+1\] Now the average velocity is the average rate of change of the position function over the time interval, so you would compute \[\frac{s(8)-s(0)}{8-0}=\frac{\left(\dfrac{1}{3}8^3-\dfrac{9}{2}8^2+18(8)+1\right)-\left(\dfrac{1}{3}0^3-\dfrac{9}{2}0^2+18(0)+1\right)}{8}=\cdots\]

OpenStudy (anonymous):

average velocity=27/8=3.3 @SithsAndGiggles please help me with question 2 and 3 also!